# Vision Mamba for Classification of Breast Ultrasound Images

### v1

@@ Abstract
Mamba-based models, VMamba and Vim, are a recent family of vision encoders that offer promising performance improvements in many computer vision tasks. This paper compares Mamba-based models with traditional Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) using the breast ultrasound BUSI and B datasets. Our evaluation, which includes multiple runs of experiments and statistical significance analysis, demonstrates that Mamba-based architectures frequently outperform CNN and ViT models with statistically significant results. These Mamba-based models effectively capture long-range dependencies while maintaining inductive biases, making them suitable for applications with limited data.
@@ 1 Introduction
Processing breast ultrasound images using deep learning can benefit from pretrained weights and a transfer learning paradigm, especially when dealing with smaller datasets. Several factors influence the effectiveness of transfer learning. One crucial factor is the choice of the pretrained dataset, and another critical factor is the selection of the encoder architecture.
@@
In this work, we adopt these architectures from natural image processing, utilizing pretrained weights trained on ImageNet [11], and compare them with common CNN and ViT architectures on two breast ultrasound datasets, BUSI [12] and B [13]. We also conduct a statistical analysis to ensure that the performance improvements are significant. Our results show that Mamba-based architectures consistently outperform other encoders with significant results. They are also never outperformed by the other encoders.
@@ 4.2 VMamba
VMamba introduces “2D Selective Scan” to the Mamba’s original scan. Furthermore, instead of breaking the image into tokens and processing each token similarly to ViT and Vim, the image patches are treated as feature maps and are processed using VSS blocks similar to CNN models, where the feature maps are down-sampled at each layer. The overall pipeline is demonstrated in Fig. 2
@@ 5 Experiments
This section summarizes the performance of various encoder types on different datasets. We use the BUSI [12] dataset (three classes), the B [13] dataset (two classes), and a combined BUSI+B dataset (three classes). Each table reports the Area Under the Curve (AUC), accuracy (ACC) metrics, and the number of parameters for each encoder. For BUSI and BUSI+B, which have three classes, the AUC is calculated individually for each class by comparing it against the other two classes. The final AUC for an experiment is then determined by averaging these individual AUC scores.
We use ResNet50 [15] and VGG16 [16] as the CNN-based models. For ViT models, we use ViT-ti16, ViT-s16, ViT-s32, ViT-b16, and ViT-b32 [2], where the prefixes ’ti,’ ’s,’ and ’b’ denote tiny, small, and base model sizes, respectively, and the numbers indicate the patch size used (16x16 or 32x32 pixels). For Mamba-based models, we use Vim-s16 [7], VMamba-ti, VMamba-s, and VMamba-b [8]. We use pretrained weights of each of these encoders from ImageNet for transfer learning.
We conduct each experiment across five training runs, using different seeds to minimize randomness. The data is split into training, validation, and test sets with respective splits of 0.7, 0.15, and 0.15. The validation set is used for early stopping; we select the checkpoint that shows the best performance on the validation set to prevent overfitting to the test set. Results are reported based on the test set’s performance for each fold.
@@ 5.1 Results on the BUSI+B Dataset
The classification performance of Mamba-based and ViT/CNN models is provided in Table 1. Mamba-based models achieve higher or comparable results to ViT and CNN models. However, the high variance in the results and the small dataset size necessitate statistical significance analysis to compare the results.
Table 1: Transfer Learning Results for BUSI+B datasets. The values of AUC and ACC are scaled between 0 to 100.
@@
VSSM VMamba-ti 29.9 $\mathbf{95.71\pm 1.01}$ $\mathbf{89.36\pm 2.33}$
@@
This analysis reveals that VMamba-ti outperforms VGG16, ViT-ti16, ViT-s32, and ViT-b32, and the difference is significant. VMamba-b outperforms VGG16 and ViT-ti16, and ViT-b16 beats VGG16 and ViT-ti16. There are no other significant differences between the models. This showcases that Mamba-based models perform similarly or outperform other encoders on the BUSI+B dataset. These results emphasize the robustness of Mamba-based models, enhancing predictive accuracy and outperforming traditional ViT and CNN models within the BUSI+B dataset.
@@ 5.2 Results on the BUSI Dataset
The classification results for the BUSI dataset are provided in Table 2. The results again suggest that the Mamba-based models outperform the other encoders on the B dataset. We ran our experiments five times and averaged the results to obtain mean and standard deviation measurements.
Table 2: Transfer Learning Results for BUSI dataset. The values of AUC and ACC are scaled between 0 to 100.
@@
VSSM VMamba-ti 29.9 $\mathbf{95.28\pm 1.89}$ $\mathbf{88.55\pm 1.67}$
@@
VMamba-b 87.5 $\mathbf{94.67\pm 2.53}$ $\mathbf{89.06\pm 3.72}$
Statistical Significance: We performed similar $p$-value tests on the BUSI dataset. The analysis demonstrates that ViT-b16, VMamba-s, and Vim-s outperform ViT-s32. VMamba-ti model outperforms ResNet50, VGG16, and ViT-s32. The VMamba-b model notably surpasses ResNet50, VGG16, ViT-ti16, ViT-s16, ViT-s32, and ViT-b32. There are no instances where Mamba-based models were outperformed by other encoders. This indicates that the Mamba-based models consistently performed at or above the level of other models on the evaluated dataset.
@@ 5.3 Results on the B Dataset
The classification results for the B dataset are provided in Table 3. VMamba-Ti is the best model, outperforming all other encoders on AUC and having a large margin on ACC. Other Mamba encoders also perform competitively and often surpass other models.
Table 3: Transfer Learning Results for B dataset. The values of AUC and ACC are scaled between 0 to 100.
@@
VSSM VMamba-ti 29.9 $\mathbf{92.66\pm 9.07}$ $\mathbf{87.50\pm 12.08}$
@@
Statistical Significance: Our statistical significance analysis for the B dataset reveals that ViT-s16 outperforms ViT-b16. Vim-s demonstrates superior performance over both ViT-ti16 and ViT-b16. VMamba-ti model shows remarkable performance by outperforming ResNet50, VGG16, ViT-ti16, ViT-s32, ViT-b16, ViT-b32, and even VMamba-b. These results underscore VMamba-ti as the top performer among the models evaluated. There are no instances where non-Mamba-based models outperformed Mamba-based models. This shows that Mamba-based models consistently achieved performances on par with or superior to those of other models evaluated in the dataset. This can be attributed to the state space nature of the Mamba methods, which allows them to capture both long- and short-distance relationships in the image.
@@ 6 Discussion
The overall performance on the three datasets demonstrates the clear advantage of Mamba-based models for breast cancer ultrasound datasets. Additionally, Mamba-based models’ ability to capture long-range dependencies while retaining some inductive bias makes them a suitable alternative to CNN and ViT models, especially in scenarios with limited data and resources. Furthermore, VMamba showed better performance overall compared to Vim. This could be due to 2D selective scan of VMamba, which allows better representation learning.
@@ 7 Conclusion
In this work, we conducted a comprehensive comparison of three families of vision encoders—ViT, CNN, and Mamba-based models—using the BUSI and B datasets and a combined BUSI+B dataset. Our evaluation included multiple runs, averages, and standard deviations, and we performed a statistical significance analysis for each experiment. Overall, Mamba-based models consistently demonstrated superior performance compared to the other models, with statistically significant results. Additionally, they were never outperformed by any other encoder type with statistically significant results.
@@ 7.0.1 Acknowledgements
This project was funded by the Natural Sciences and Engineering Research Council of Canada (NSERC) and Fonds de recherche du Québec – Nature et technologies (FRQNT).

### v2

@@ Abstract
Mamba-based models, VMamba and Vim, are a recent family of vision encoders that offer promising performance improvements in many computer vision tasks. This paper compares Mamba-based models with traditional Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) using the breast ultrasound BUSI dataset and Breast Ultrasound B dataset. Our evaluation, which includes multiple runs of experiments and statistical significance analysis, demonstrates that some of the Mamba-based architectures often outperform CNN and ViT models with statistically significant results. For example, in the B dataset, the best Mamba-based models have a 1.98% average AUC and a 5.0% average Accuracy improvement compared to the best non-Mamba-based model in this study. These Mamba-based models effectively capture long-range dependencies while maintaining some inductive biases, making them suitable for applications with limited data. The code is available at https://github.com/anasiri/BU-Mamba.
@@ 1 Introduction
Processing breast ultrasound images using deep learning can benefit from pretrained weights and a transfer learning paradigm, especially when dealing with smaller datasets [1]. Several factors influence the effectiveness of transfer learning. One crucial factor is the choice of the pretrained dataset [2], and another critical factor is the selection of the encoder architecture [2].
@@
In this work, we adopt these architectures from natural image processing, utilizing pretrained weights trained on ImageNet [13], and compare them with common CNN and ViT architectures on two breast ultrasound datasets, BUSI [14] and B dataset[15]. We also conduct a statistical analysis to ensure that the performance improvements are significant. Our results show that Mamba-based architectures frequently outperform other encoder types with significant results. They are also never outperformed by the other encoders in terms of statistical significance.
@@ 4.2 VMamba
VMamba introduces “2D Selective Scan” to the Mamba’s original scan. Furthermore, instead of breaking the image into tokens and processing each token similarly to ViT and Vim, the image patches are treated as feature maps and are processed using VSS blocks similar to CNN models, where the feature maps are down-sampled at each layer. The overall pipeline is demonstrated in Fig. 2.
@@ 5 Experiments
This section summarizes the performance of various encoder types on different datasets. For our classification task, we use the BUSI [14] dataset (three classes of benign, malignant, and normal), the Dataset B of [15] (two classes of benign and malignant), and a combined BUSI+B dataset (three classes of normal, benign, and malignant). Each table reports the Area Under the Curve (AUC), Accuracy (ACC) metrics, and the number of parameters for each encoder. Since the experiment datasets are imbalanced, AUC is required as a more valuable metric. For BUSI and BUSI+B, which have three classes, the AUC is calculated individually for each class by comparing it against the other two classes. The final AUC for an experiment is then determined by averaging these individual AUC scores.
We use ResNet50 [17] and VGG16 [18] as the CNN-based models. For ViT models, we use ViT-ti16, ViT-s16, ViT-s32, ViT-b16, and ViT-b32 [4], where the prefixes ’ti,’ ’s,’ and ’b’ denote tiny, small, and base model sizes, respectively, and the numbers indicate the patch size used (16x16 or 32x32 pixels). For Mamba-based models, we use Vim-s16 [9], VMamba-ti, VMamba-s, and VMamba-b [10]. These Mamba-based models have the default parameters and definition as provided in the papers. We use pretrained weights of each of these encoders from ImageNet for transfer learning.
We conduct each experiment across five training runs, using different seeds to minimize randomness in experiments. The data is split into training, validation, and test sets with respective splits of 0.7, 0.15, and 0.15. The validation set is used for early stopping; we select the checkpoint that shows the best performance on the validation set to prevent overfitting to the test set. Results are reported based on the test set’s performance for each fold.
@@ 5.1 Results on the BUSI+B Dataset
The classification performance of Mamba-based and ViT/CNN models is provided in Table 1. We ran our experiments five times and averaged the results to obtain mean and standard deviation measurements. Mamba-based models achieve higher or comparable results to ViT and CNN models. However, the high variance in the results and the small dataset size necessitate statistical significance analysis to compare the results.
Table 1: Transfer Learning Results for BUSI+B datasets. The values of AUC and ACC are scaled between 0 to 100. The results are averaged over five runs.
@@
VSSM VMamba-ti 29.9 $95.71\pm 1.01$ $89.36\pm 2.33$
@@
This analysis reveals that VMamba-ti outperforms VGG16 (p-value: 0.004), ViT-ti16 (p-value: 0.003), ViT-s32 (p-value: 0.014), and ViT-b32 (p-value: 0.022), and the difference is significant. VMamba-b outperforms VGG16 (p-value: 0.037) and ViT-ti16 (p-value: 0.015). Additionally, ViT-b16 performs better than VGG16 (p-value: 0.044) and ViT-ti16 (p-value: 0.013). There are no other significant differences between pairs of models (the p-value is bigger than 0.05). This showcases that Mamba-based models perform similarly or outperform other encoders on the BUSI+B dataset with respect to statistical significance.
It is important to note that while ResNet50, as shown in Table 1, achieves a higher mean AUC of $95.74$ compared to some of the other models, the statistical significance analysis indicates that this improvement is not significant. For example, when testing for ResNet50 against ViT-s32, the p-value is 0.27, even though the mean AUC and ACC for ResNet50 are better. Due to the experiments’ small dataset size and inherent randomness, we report both mean and standard deviation for each metric, which makes direct comparisons between models more complex. Therefore, statistical analysis is necessary to reduce this randomness and better identify instances where the differences between pairs of models are statistically significant and not due to randomness in the experiments.
@@ 5.2 Results on the BUSI Dataset
The classification results for the BUSI dataset are provided in Table 2. We ran our experiments five times and averaged the results to obtain the mean and standard deviation measurements. The results suggest that the Mamba-based models outperform the other encoders on the B dataset.
Table 2: Transfer Learning Results for BUSI dataset. The values of AUC and ACC are scaled between 0 to 100. The results are averaged over five runs.
@@
VSSM VMamba-ti 29.9 $95.28\pm 1.89$ $88.55\pm 1.67$
@@
VMamba-b 87.5 $94.67\pm 2.53$ $89.06\pm 3.72$
Statistical Significance: We performed $p$-value tests on the BUSI dataset. The analysis demonstrates that ViT-b16, VMamba-s, and Vim-s outperform ViT-s32 (with a p-value of 0.036, 0.027, and 0.011, respectively). VMamba-ti model outperforms ResNet50 (p-value: 0.032), VGG16 (p-value: 0.024), and ViT-s32 (p-value: 0.002). The VMamba-b model notably surpasses ResNet50 (p-value: 0.015), VGG16 (p-value: 0.006), ViT-ti16 (p-value: 0.018), ViT-s16 (p-value: 0.022), ViT-s32 (p-value: 0.0001), and ViT-b32 (p-value: 0.029). There are no other pairs with p-values smaller than 0.05. These results indicate that the Mamba-based models frequently performed at or above the level of other models with respect to statistical significance.
@@ 5.3 Results on the B Dataset
The classification results for the B dataset are provided in Table 3. We conducted our experiments five times and calculated the average and standard deviation of the results. VMamba-ti is the best model, outperforming all other encoders on average AUC and having a large margin on average ACC. Other Mamba encoders also perform competitively and often surpass other models.
Table 3: Transfer Learning Results for B dataset. The values of AUC and ACC are scaled between 0 to 100. The results are averaged over five runs.
@@
VSSM VMamba-ti 29.9 $92.66\pm 9.07$ $87.50\pm 12.08$
@@
Statistical Significance: Our statistical significance analysis for the B dataset reveals that ViT-s16 outperforms ViT-b16 (p-value: 0.034). Vim-s performs better than both ViT-ti16 (p-value: 0.032) and ViT-b16 (p-value: 0.019). VMamba-ti model shows remarkable performance by outperforming ResNet50 (p-value: 0.011), VGG16 (p-value: 0.028), ViT-ti16 (p-value: 0.004), ViT-s32 (p-value: 0.012), ViT-b16 (p-value: 0.003), ViT-b32 (p-value: 0.007), and even VMamba-b (p-value: 0.034). There are no other pairs of models with p-values smaller than 0.05. These results underscore VMamba-ti as the top performer among the models evaluated. Furthermore, there are no instances with statistical significance where non-Mamba-based models outperformed Mamba-based models.
@@ 6 Discussion
The overall performance on the three datasets demonstrates the advantages of using Mamba-based models for breast cancer ultrasound datasets. Additionally, Mamba-based models’ ability to capture long-range dependencies while retaining some inductive bias makes them a suitable alternative to CNN and ViT models, especially in scenarios with limited data and resources. Furthermore, VMamba showed better performance overall compared to Vim. This could be due to 2D selective scan of VMamba, which allows better representation learning.
One limitation of this study is that the statistical significance analysis is not considering the imbalance in the dataset due the multi-class analysis and complexity of comparing AUC for pairs of models. Despite this, some Mamba-based models frequently have higher average AUC as provided in the experiments section. A more comprehensive statistical analysis, including the multi-class AUC, would provide stronger support for our results.
@@ 7 Conclusion
In this work, we conducted a comprehensive comparison of three families of vision encoders—ViT, CNN, and Mamba-based models—using the BUSI and B datasets and a combined BUSI+B dataset. Our evaluation included multiple runs, averages, and standard deviations, and we performed a statistical significance analysis for each experiment. Overall, Mamba-based models frequently demonstrated competitive performance compared to the other models, with statistically significant results. Additionally, they were never outperformed by any other encoder type regarding statistical significance.
@@ 7.0.1 Acknowledgements
This project was funded by the Natural Sciences and Engineering Research Council of Canada (NSERC) and Fonds de recherche du Québec – Nature et technologies (FRQNT). We also extend our gratitude to Behnaz Gheflati for their valuable time.